# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ 5.1 Preparations
Let $\bm{Z},\bm{W}\in\mathbb{R}^{m\times N}$ be random. For any probability measure $\mu\in\mathcal{P}(\mathbb{R}^{m})$,
$\Big{|}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},\mu)}\Big{|}\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}\,,$
$0\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}-\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{Z}),\mathrm{Law}(\bm{W}))\leq 2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}\,.$
Proof of Lemma 5.1.
By the triangle inequality,
$|\mathcal{W}_{2}(q_{\bm{Z}},q_{\bm{W}})-\mathcal{W}_{2}(q_{\bm{Z}},\mu)|\leq\mathcal{W}_{2}(q_{\bm{W}},\mu)\,.$
By Minkowski’s inequality,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}\geq\sqrt{\mathbb{E}|\mathcal{W}_{2}(q_{\bm{Z}},q_{\bm{W}})-\mathcal{W}_{2}(q_{\bm{Z}},\mu)|^{2}}\geq\Big{|}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},\mu)}\Big{|}\,.$
Therefore, for any coupling $\gamma$ between $\mathrm{Law}(\bm{Z})$ and $\mathrm{Law}(\bm{W}))$, the quantity $\sqrt{\mathbb{E}_{(\bm{Z}^{\prime},\bm{W}^{\prime})\sim\gamma}\mathcal{W}_{2}^{2}(q_{\bm{Z}^{\prime}},q_{\bm{W}^{\prime}})}$ belongs to a fixed interval of length $2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{W}},\mu)}$. This immediately proves the desired result. ∎
5.2 Main proof
Let $\bm{Y}\in\mathbb{R}^{m\times N}$ be a random array such that the $N$ particles $\{\bm{Y}^{:,j}\}_{j=1}^{N}$ are i.i.d. samples from $q_{*}$. When $N$ is large, $q_{\bm{Y}}$ concentrates around $q_{*}$ (Corollary 6.1).
In addition, let $\{B_{t}^{i,j}\}_{t\geq 0}$ for $(i,j)\in[m]\times[N]$ be i.i.d. standard Brownian motions. Suppose that $\bm{Y}$ and $\{\bm{B}_{t}\}_{t\geq 0}$ are independent, and both are independent of the randomness in Algorithm 2. For $(i,j)\in[m]\times[N]$, define $\{Y_{t}^{i,j}\}_{t\geq 0}$ as the unique strong solution to the stochastic differential equation
$\begin{cases}&\mathrm{d}Y_{t}^{i,j}=-\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i,j}\\ &Y_{0}^{i,j}=Y^{i,j}\,.\end{cases}$
We present a useful lemma on the regularity of $\{\bar{V}_{i}\}_{i=1}^{m}$, which guarantees the well-posedness of the above SDE (Karatzas and Shreve, 1991). The proof is immediate and thus omitted.
Lemma 5.2.
Let Assumption 4.1 and 4.2 hold. For any $\mu\in\mathcal{P}(\mathbb{R}^{m-1})$ and $x\in\mathbb{R}$, we have $\alpha\leq\bar{V}_{i}^{\prime\prime}(x,\mu)\leq L$ and $|V_{i}^{\prime\prime\prime}(x,\mu)|\leq\widetilde{L}$.
By definition, we have
$Y_{h}^{i,j}=Y_{0}^{i,j}-\int_{0}^{h}\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}B_{h}^{i,j}\approx Y_{0}^{i,j}-h\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2h}(B_{h}^{i,j}/\sqrt{h})$
and $B_{h}^{i,j}/\sqrt{h}\sim N(0,1)$. Hence, the update from $\bm{Y}_{0}$ to $\bm{Y}_{h}$ is similar to that from $\bm{X}_{n}$ to $\bm{X}_{n+1}$. This observation allows us to bound $\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n+1}),\mathrm{Law}(\bm{Y}_{h}))$ using $\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n}),\mathrm{Law}(\bm{Y}_{0}))$. For any $t\geq 0$, $\bm{Y}_{t}$ has the same distribution as $\bm{Y}$. We can establish the following recursive inequality. See Section 5.3 for its proof.
Lemma 5.3.
Let Assumptions 4.1 and 4.2 hold. Suppose that $0<h<\min\{2/(\alpha+L),B\alpha/(4L^{2})\}$. Define $W_{n}=\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n}),\mathrm{Law}(\bm{Y}))$ for $n\in\mathbb{N}$. We have
$W_{n+1}^{2}\leq\bigg{(}(1-\alpha h/4)W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}+4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha).$
Next, we invoke a useful lemma.
Lemma 5.4 (Lemma 1 in Dalalyan and Karagulyan (2019)).
Let $A$, $B$ and $C$ be non-negative numbers such that $A\in(0,1)$. Assume that a sequence of non-negative numbers $\{x_{n}\}_{n=0}^{\infty}$ satisfies the recursive inequality
$x_{n+1}^{2}\leq[(1-A)x_{n}+C]^{2}+B^{2},\qquad\forall n\in\mathbb{Z}_{+}.$
Then,
$x_{n}\leq(1-A)^{n}x_{0}+\frac{C}{A}+\frac{B}{\sqrt{A}},\qquad\forall n\in\mathbb{Z}_{+}.$
Applying Lemma 5.4 to the sequence $\{W_{n}\}_{n=0}^{\infty}$ in Lemma 5.3, we get
$W_{n}$ $\leq(1-\alpha h/4)^{n}W_{0}+\frac{2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}}{\alpha h/4}+\sqrt{\frac{4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha)}{\alpha h/4}}$
$\leq(1-\alpha h/4)^{n}W_{0}+\frac{8\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}}{\alpha h}+4\sqrt{mh\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{3}+\widetilde{L}^{2}/\alpha^{2})}.$
By Lemma 5.1,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}$
$\leq\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{n}),\mathrm{Law}(\bm{Y}))+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}$
$=W_{n}+3\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}.$
On the other hand,
$W_{0}=\widetilde{\mathcal{W}}(\mathrm{Law}(\bm{X}_{0}),\mathrm{Law}(\bm{Y}))\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{\bm{Y}})}\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}.$
Based on the above estimates, the assumption $h<2/(\alpha+L)$, and Corollary 6.1, there exists a universal constant $C$ such that
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}$
$\quad+\frac{C\sqrt{m}}{\alpha}\bigg{(}h^{-1}\sqrt{\frac{\log N}{\alpha N}}+\sqrt{h\max\{h,(LB)^{-1}\}(L^{4}/\alpha+\widetilde{L}^{2})}\bigg{)}.$
It remains to prove Lemma 5.3.
5.3 Proof of Lemma 5.3
5.3.1 Step 1: Analysis of $W_{n}$
According to Lemma 5.1,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})}\leq W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}.$ (5.1)
We now look into the quantity $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})$. Let $X_{n}^{i,(1)}\leq\cdots\leq X_{n}^{i,(N)}$ and $Y^{i,(1)}\leq\cdots\leq Y^{i,(N)}$ be the order statistics of $\{X_{n}^{i,j}\}_{j=1}^{N}$ and $\{Y^{i,j}\}_{j=1}^{N}$, respectively. Then,
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|X_{n}^{i,(j)}-Y^{i,(j)}|^{2},\qquad\forall i\in[m].$
Denote by $S_{N}$ the set of all permutations of $[N]$. There exist $\{\tau_{i}\}_{i=1}^{m}\subseteq S_{N}$, where $\tau_{i}$ is determined by $\bm{X}_{n}^{i,:}$ and $\bm{Y}^{i,:}$, such that
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|X_{n}^{i,\tau_{i}(j)}-Y^{i,j}|^{2},\qquad\forall i\in[m].$
For each $i\in[m]$, $\{Y^{i,j}\}_{j=1}^{N}$ are i.i.d. from an absolutely continuous distribution $q_{*}^{i}$. The continuity of $q_{*}^{i}$ and the symmetry of $\{Y^{i,j}\}_{j=1}^{N}$ imply the following.
Fact 5.1.
Conditioned on $\bm{X}_{n}$, the permutations $\{\tau_{i}\}_{i=1}^{m}$ are
  • uniquely defined with probability 1;
  • independent and uniformly distributed on $S_{N}$.
Define an array $\bm{U}$ with $U^{i,j}=X_{n}^{i,\tau_{i}(j)}$. Fact 5.1 implies that given $\bm{X}_{n}$, $\{\bm{U}^{:,j}\}_{j=1}^{N}$ are i.i.d. samples from $q_{\bm{X}_{n}}$. In addition, we have $q_{\bm{U}}=q_{\bm{X}_{n}}$ and
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|U^{i,j}-Y^{i,j}|^{2},\qquad\forall i\in[m].$
By Lemma 6.1,
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}}^{i})=\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}|U^{i,j}-Y^{i,j}|^{2}.$ (5.2)
5.3.2 Step 2: Analysis of $W_{n+1}$
Define two arrays $\bm{M},\bm{W}\in\mathbb{R}^{m\times N}$ through
$M^{i,j}=U^{i,j}-hg_{i}(U^{i,j})+\sqrt{2}B_{h}^{i,j},$
$W^{i,j}=U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})+\sqrt{2}B_{h}^{i,j}.$
It is easily seen that $q_{\bm{M}}$ and $q_{\bm{X}_{n+1}}$ have the same distribution, and $\bm{W}=\mathbb{E}(\bm{M}|\bm{X}_{n},\{\bm{Y}_{t}\}_{t=0}^{h},\{\bm{B}_{t}\}_{t=0}^{h})$.
For any $t\geq 0$, $\bm{Y}_{t}$ has the same distribution as $\bm{Y}$. Hence,
$W_{n+1}^{2}=\widetilde{\mathcal{W}}^{2}(\mathrm{Law}(\bm{X}_{n+1}),\mathrm{Law}(\bm{Y}))\leq\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{M}},q_{\bm{Y}_{h}})\leq\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|M^{i,j}-Y_{h}^{i,j}|^{2}.$
By $\bm{W}=\mathbb{E}(\bm{M}|\bm{X}_{n},\{\bm{Y}_{t}\}_{t=0}^{h},\{\bm{B}_{t}\}_{t=0}^{h})$, we have
$\mathbb{E}|M^{i,j}-Y_{h}^{i,j}|^{2}=\mathbb{E}[(M^{i,j}-W^{i,j})+(W^{i,j}-Y_{h}^{i,j})]^{2}=\mathbb{E}|M^{i,j}-W^{i,j}|^{2}+\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}.$
Hence,
$W_{n+1}^{2}\leq\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\Big{(}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}+\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}\Big{)}.$ (5.3)
Below we look into the quantities on the right-hand side.
5.3.3 Step 3: Analysis of $\mathbb{E}|M^{i,j}-W^{i,j}|^{2}$

### v3

@@ 5 Proofs
Suppose that Assumption 4.1 holds and $0<h<\min\{\frac{2}{\alpha+L},\frac{B\alpha}{4L^{2}}\}$. There exists a stationary process $\{\bm{Y}_{t}\}_{t\geq 0}$ with $\bm{Y}_{t}\in\mathbb{R}^{m\times N}$, such that for any $t\geq 0$, $\{\bm{Y}_{t}^{:,j}\}_{j=1}^{N}$ are i.i.d. samples from $q_{*}$, and
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n+1}},q_{\bm{Y}_{(n+1)h}})\leq(1-\alpha h/2)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})+\frac{4mh^{2}L^{2}}{\alpha}\bigg(1+\frac{1}{2B}+\frac{hL^{2}}{\alpha}\bigg).$
Define $W_{n}=\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})$. When $0<h<\alpha/(4L^{2})$, Lemma˜5.1 implies that $W_{n+1}\leq(1-A)W_{n}+R$, where $A=\alpha h/2$ and $R=8mh^{2}L^{2}/\alpha$. By induction, we get
$W_{n}\leq(1-A)^{n}W_{0}+\sum_{k=0}^{n-1}(1-A)^{k}R\leq(1-A)^{n}W_{0}+\frac{R}{A}$
and thus $\sqrt{W_{n}}\leq(1-A)^{n/2}\sqrt{W_{0}}+\sqrt{R/A}$. Then,
$\sqrt{W_{n}}\leq(1-A)^{n/2}\sqrt{W_{0}}+\sqrt{R/A}=(1-\alpha h/2)^{n/2}\sqrt{W_{0}}+\frac{4L}{\alpha}\sqrt{mh}.$ (5.1)
We now relate $W_{n}$ to $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})$. By Lemma˜6.1, we have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})}\geq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{nh}},q_{*})},$
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{\bm{Y}_{0}})}\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}},q_{*})}.$
The stationarity of $\{\bm{Y}_{t}\}_{t\geq 0}$ implies $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{nh}},q_{*})=\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}},q_{*})$. By the above estimates and (5.1),
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/2)^{n/2}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\frac{4L}{\alpha}\sqrt{mh}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}},q_{*})}.$ (5.2)
Note that the columns of $\bm{Y}_{0}$ are i.i.d. samples from $q_{*}$. By Lemma˜6.4, each $q_{*}^{i}$ is log-concave with variance bounded by $\alpha^{-1}$. Then, Corollary˜6.1 implies that $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}},q_{*})=O(\frac{m\log N}{\alpha N})$. This completes the proof of Theorem˜4.1. It remains to prove Lemma˜5.1, which can be found in Section˜5.5.
5.2 Proof of the claim regarding the particle approximation error
Note that for $i\in[m]$, $q_{\bm{Y}}^{i}$ is an empirical measure formed by $N$ independent samples from the marginal distribution $q^{i}=N(0,\alpha^{-1})$. Corollary 6.14 in BLe19 shows the existence of a universal constant $c$ such that $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}}^{i},q^{i})\geq\frac{c\log\log N}{\alpha N}$ holds for all $N\geq 3$. By Lemma˜6.2, we have
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q)=\sum_{i=1}^{m}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}}^{i},q^{i})\geq\frac{cm\log\log N}{\alpha N}.$
5.3 Proof of Theorem˜4.2
Examining the proof of Theorem˜4.1, we see that the derivation of (5.2) works seamlessly for the block case. The result now reads
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/2)^{n/2}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\frac{4L}{\alpha}\sqrt{dh}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}},q_{*})},$
where $\bm{Y}_{0}\in\mathbb{R}^{d\times N}$ is a random array whose columns are i.i.d. samples from $q_{*}$, and $q_{\bm{Y}_{0}}$ is constructed according to the block version of Definition˜3.2. More concretely, let $Y_{0}^{i,j}$ be the entries of $\bm{Y}_{0}$ in the $(1+\sum_{k=1}^{i-1}d_{k})$-th through $\sum_{k=1}^{i}d_{k}$-th rows and the $j$-th column. Then, $q_{\bm{Y}_{0}}$ is defined by $q_{\bm{Y}_{0}}^{1}\otimes\cdots\otimes q_{\bm{Y}_{0}}^{m}$ with $q_{\bm{Y}_{0}}^{i}=\frac{1}{N}\sum_{j=1}^{N}\delta_{Y_{0}^{i,j}}$. By Lemma˜6.2, $\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}},q_{*})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{Y}_{0}}^{i},q_{*}^{i})$. We complete the proof using the fact that $q_{\bm{Y}_{0}}^{i}$ is an empirical distribution formed by $N$ independent samples from $q_{*}^{i}$.
5.4 Proof of Corollary˜4.1
When $N\geq 9$, we have $N^{-1}\log N<1/4$ and $h=\frac{\alpha\log N}{L^{2}N}<\alpha/(4L^{2})$. Hence, Theorem˜4.1 holds. Since $1-x\leq e^{-x}$ for all $x\in\mathbb{R}$, we have $(1-\alpha h/2)^{n/2}\leq e^{-\alpha hn/4}$ and
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq\exp\bigg(-\frac{n\alpha^{2}\log N}{4L^{2}N}\bigg)\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+2C\sqrt{\frac{m\log N}{\alpha N}}.$
By Lemma˜6.1, we have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{0})}+\mathcal{W}_{2}(q_{0},q_{*}).$
Corollary˜6.1 guarantees $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{0})=O(\frac{m\log N}{\alpha N})$. Then, we obtain the desired result by redefining $C$.
5.5 Proof of Lemma˜5.1
The proof is based on a coupling argument.
5.5.1 Step 1: Construction of coupling
To begin with, let $\bm{X}_{0}\in\mathbb{R}^{m\times N}$ be the initial value of Algorithm 2 that is allowed to be random; $\{B_{t}^{i,j}\}_{t\geq 0}$, $(i,j)\in[m]\times[N]$ be i.i.d. standard Brownian motions that are independent of $\bm{X}_{0}$. By running Algorithm 2 with $\bm{\xi}_{n}^{i,:}$ replaced by $(\bm{B}_{(n+1)h}^{i,:}-\bm{B}_{nh}^{i,:})/\sqrt{h}$, we obtain a sequence $\{\bm{X}_{n}\}_{n=0}^{\infty}$ that has the same distribution as the one in Algorithm 2.
Next, we construct an auxiliary process $\{\bm{Y}_{t}\}_{t\geq 0}$ that is coupled with $\{\bm{X}_{n}\}_{n=0}^{\infty}$. Let $\bm{Y}_{0}\sim q_{*}$ be independent of the above construction of $\{\bm{X}_{n}\}_{n=0}^{\infty}$. Choose any $n\in\mathbb{N}$ and suppose that $\{\bm{Y}_{t}\}_{0\leq t\leq nh}$ has been defined. We now define $\{\bm{Y}_{t}\}_{nh<t\leq(n+1)h}$. For each $i\in[m]$, the Birkhoff–von Neumann algorithm implies the existence of a permutation $\tau_{n,i}$ of $[N]$, determined by $\bm{X}_{n}^{i,:}$ and $\bm{Y}_{nh}^{i,:}$, such that
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}_{nh}}^{i})=\frac{1}{N}\sum_{j=1}^{N}|X_{n}^{i,j}-Y_{nh}^{i,\tau_{n,i}(j)}|^{2}.$ (5.3)
See Chapter 2 of Vil09. Define $\{\bm{Y}_{t}\}_{nh\leq t\leq(n+1)h}$ as the strong solution to the SDE
$\mathrm{d}Y_{t}^{i,j}=-\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i,\tau_{n,i}^{-1}(j)},\qquad nh\leq t\leq(n+1)h.$ (5.4)
Let $\bar{B}_{0}^{i,j}=0$ for all $(i,j)\in[m]\times[N]$. For every $n\in\mathbb{N}$ and $t\in(nh,(n+1)h]$, define $\bar{B}_{t}^{i,j}=\bar{B}_{nh}^{i,j}+(B_{t}^{i,\tau_{n,i}^{-1}(j)}-B_{nh}^{i,\tau_{n,i}^{-1}(j)})$. Then, $\{\bar{B}_{t}^{i,j}\}_{t\geq 0}$, $(i,j)\in[m]\times[N]$ are i.i.d. standard Brownian motions that are independent of $\bm{Y}_{0}$. Our process $\{\bm{Y}_{t}\}_{t\geq 0}$ solves the SDE
$\mathrm{d}Y_{t}^{i,j}=-\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}\mathrm{d}\bar{B}_{t}^{i,j}.$
Hence, for any $t\geq 0$, $\{\bm{Y}_{t}^{:,j}\}_{j=1}^{N}$ are i.i.d. samples from $q_{*}$. Below we will analyze $\{\bm{X}_{n}\}_{n=0}^{\infty}$ and $\{\bm{Y}_{t}\}_{t\geq 0}$.
5.5.2 Step 2: Error decomposition
Choose any $n\in\mathbb{N}$. By (5.3) and Lemma˜6.2,
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}}^{i},q_{\bm{Y}_{nh}}^{i})=\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}|X_{n}^{i,j}-Y_{nh}^{i,\tau_{n,i}(j)}|^{2}.$ (5.5)
Meanwhile, Lemma˜6.2 and the definition of $\mathcal{W}_{2}$ imply that
$\mathcal{W}_{2}^{2}(q_{\bm{X}_{n+1}},q_{\bm{Y}_{(n+1)h}})$ $=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n+1}}^{i},q_{\bm{Y}_{(n+1)h}}^{i})$
$\leq\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}|X_{n+1}^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}.$ (5.6)
We will show how the error evolves after one iteration. Note that
$X_{n+1}^{i,j}=X_{n}^{i,j}-hg_{i}(X_{n}^{i,j})+\sqrt{2}(B_{(n+1)h}^{i,j}-B_{nh}^{i,j}).$
Define another array $\bm{W}\in\mathbb{R}^{m\times N}$ through
$W^{i,j}=X_{n}^{i,j}-h\bar{V}_{i}^{\prime}(X_{n}^{i,j},q_{\bm{X}_{n}}^{-i})+\sqrt{2}(B_{(n+1)h}^{i,j}-B_{nh}^{i,j}).$
We have $\bm{W}=\mathbb{E}(\bm{X}_{n+1}|\bm{X}_{n},\{\bm{B}_{t}\}_{nh\leq t\leq(n+1)h},\{\bm{Y}_{t}\}_{nh\leq t\leq(n+1)h})$. It can be viewed as the update by Algorithm 1 with full gradient. Then,
$\mathbb{E}|X_{n+1}^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}$ $=\mathbb{E}[(X_{n+1}^{i,j}-W^{i,j})+(W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)})]^{2}$
$=\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}+\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}.$
From this and (5.6), we obtain that
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n+1}},q_{\bm{Y}_{(n+1)h}})\leq\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\Big(\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}+\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}\Big).$ (5.7)
Below, we look into the quantities on the right-hand side. The first term $\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}$ is the error incurred by stochastic gradient approximation and hence vanishes as $B\to\infty$. The second term $\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,j}|^{2}$ bounds the error caused by finite-difference approximation.
5.5.3 Step 3: Analysis of $\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}$
We will prove that
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}\leq\frac{2h^{2}}{B}[L^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}].$ (5.8)